OpenStudy (anonymous):

Evaluate the limit using l'Hopital's Rule (I tried this, but I can't figure out where I went wrong :/) lim(y->pi/2)((pi/2)-y)

OpenStudy (anonymous):

lim(y−>π/2)(π2−y)\[\lim(y−>π/2)(\frac{ \pi }{ 2 }−y)tany\]

OpenStudy (anonymous):

ignore the first part

OpenStudy (anonymous):

@jim_thompson5910 if you can help, it would be greatly appreciated :)

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@tkhunny

OpenStudy (tkhunny):

In what state is in an Indeterminate Form?

OpenStudy (anonymous):

what

OpenStudy (tkhunny):

Not a good response. In order to utilize l'Hospital's Rule, you must have an Indeterminate Form. Do you?

OpenStudy (anonymous):

Well I would've said California, but I figured that that was the wrong answer :3 Yes we do have an indeterminate form

OpenStudy (tkhunny):

Prove it. It's not obvious.

OpenStudy (anonymous):

0 times infinity

OpenStudy (tkhunny):

That doesn't contain a Numerator and Denominator, which is where the rule is useful. Is there a firm like that?

OpenStudy (tkhunny):

* form

OpenStudy (anonymous):

Yes there is. You have to rewrite with an LCD

OpenStudy (anonymous):

Then you get a fraction

OpenStudy (tkhunny):

Show me.

OpenStudy (anonymous):

Its what my teacher told me to do

OpenStudy (tkhunny):

Okay. Show me.

OpenStudy (anonymous):

Oh wait, thats infinity-infinity

OpenStudy (anonymous):

For this one, its with natural logs

OpenStudy (anonymous):

so you take the natural log of the inside part

OpenStudy (anonymous):

Then the exponent can be rewritten as a multiplier

OpenStudy (anonymous):

Then you rewrite as a fraction

OpenStudy (tkhunny):

Are we talking about the same problem? \(\lim_{x\rightarrow \pi/2}\left(\pi/2 - y\right)\cdot \tan{y}\)

OpenStudy (anonymous):

Yes

OpenStudy (tkhunny):

Sorry, 'y', not 'x'.

OpenStudy (anonymous):

Wait i was wrong again

OpenStudy (anonymous):

Sorry

OpenStudy (tkhunny):

Then why not just \(\dfrac{\pi/2 - y}{\cot(y)}\)? This gives the form 0/0.

OpenStudy (anonymous):

This is the one where you rewrite as division

OpenStudy (tkhunny):

Don't get hung up on techniques. Get hung up on when the rule is applicable and then just get it where it is useful.

OpenStudy (anonymous):

I did\[tany/(1/((\pi/2)-y)))\]

OpenStudy (anonymous):

Which is infinity over infinity

OpenStudy (anonymous):

Then I l'hopitaled

OpenStudy (anonymous):

And got sec^2x/(-1/((pi/2)-y))^2)

OpenStudy (anonymous):

And I got hung up because thats infinity over infinity as well, and l'hopitaling again wouldn't help me in any way

OpenStudy (tkhunny):

Well, that's ONLY because you chose poorly when you moved something to the denominator. You already have a linear factor in the numerator. That will vanish with the first derivative. Don't move it. Move the tangent to the denominator.

OpenStudy (anonymous):

I didn't realize that there were rules for choosing

OpenStudy (tkhunny):

The sign was wrong, too. But you can worry about that, later.

OpenStudy (anonymous):

We only worked with x's as the multiplier. Then you could just rewrite those as 1/x

OpenStudy (tkhunny):

There are not rules for choosing. Don't get hung up on such things. You can, however, choose wisely in many circumstances.

OpenStudy (anonymous):

I just assumed to do the term with the x in it

OpenStudy (tkhunny):

Yup. It got you nowhere. Move the tangent and give it a go.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

I got -y/-csc^2y

OpenStudy (anonymous):

So thats

OpenStudy (anonymous):

pi/2/0

OpenStudy (anonymous):

so infinity

OpenStudy (tkhunny):

-y? Try that numerator, again.

OpenStudy (anonymous):

oh snap

OpenStudy (anonymous):

-1. sorry bout that

OpenStudy (anonymous):

And of course I messed up the unit circle. Its 1

OpenStudy (anonymous):

Thank you so much for your help. How do you decide which part to pick to rewrite as division?

OpenStudy (tkhunny):

It's not always obvious. With a little smirk on my face, you pick a way that is useful. :-) If you find that one way is not useful, try something else.

OpenStudy (tkhunny):

What did you get for the final answer?

OpenStudy (anonymous):

So basically the same thing as integration by parts XD

OpenStudy (anonymous):

1

OpenStudy (tkhunny):

Perfect. Good work.

OpenStudy (tkhunny):

Fair enough. Keep in mind what you will be differentiating. If you only make things more difficult, that's not likely to be helpful. Keep up the good work.

OpenStudy (anonymous):

Thank you very much!